Solve algebraically, using one variable. The length of a rectangle is three less than twice the width. If the perimeter is 24 inches, what are the length and width of the rectangle?
Answer by mxgirl22(39)


Solve algebraically, using one variable. The length of a rectangle is three less than twice the width. If the perimeter is 24 inches, what are the length and width of the rectangle?

-------------------------------------------
1.) Get key information out.

length (l) of a rectangle is(=) three (3) less (-) than twice(x2) the width(w).

So length = 2(times)width-3

1ength = 2w-3

width = w
perimeter is(=) 24 inches
perimeter = 24 inches

------------------------------------------
2.) Set up your equation.

***Note: Formula for perimeter is: 2 ( l + w )

2 ( l + w ) = 24


2[(2w-3)+(w)] = 24

(4w-6)+(2w) = 24

4w-6+2w = 24

6w-6 = 24

6w-6+6 = 24+6

6w = 30

6w/6 = 30/6

w = 5

------------------------------------------
3.) Using the value we got for "w" in step 2, plug in to what we said the length is equal to in step 1. (1 = 2w-3 ).

length= 2w-3

l=2(5)-3

l=10-3

l=7

------------------------------------------
4.) Check with perimeter formula. p = 2( l + w )

Use l=7, w=5, p=24


p = 2( l + w )

24= 2[ (7)+(5) ]

24=14+10

24=24

Answers check!

--------------------------------------
5.) Answer.


If the perimeter is 24 inches, the length of the rectangle is 7 inches and width is 5 inches.